The supplement of angle x is 60 degrees more than twice the compliment of angle x. Find the measure of angle x.
60+ 2x=180-60
2x/2= 120/2
x=60

Answer by stanbon(75887)

The supplement of angle x is 60 degrees more than twice the compliment of angle x. Find the measure of angle x.
-----------------------
The supplement of angle x is 180-x
The complement of angle x is 90-x
-----
Equation:
supplement = 2*complement + 60
180-x = 2(90-x)+60
180-x = 180-2x+60
x = 60 degrees
======================
Cheers,
Stan H.